OpenStudy (goldps):

how do i get the answer for 1/2 * sqrt(128) * sqrt(50)

OpenStudy (alekos):

first step is to simplify sqrt(128) 128 = 64x2 so sqrt(64x2) = 8sqrt2

OpenStudy (alekos):

can you see how I got that?

OpenStudy (goldps):

ya, sry, was doin other problems go on

OpenStudy (alekos):

so now we have 1/2*8*sqrt(2)*sqrt(50) = 4*sqrt(2)*sqrt(50) now we just combine the square roots

OpenStudy (alekos):

sqrt(2)*sqrt(50) = what do you think?

OpenStudy (goldps):

sqrt(2*50)=sqrt(100) = 10 then 4*10 = 40 40 = 1/2 * sqrt(128)*sqrt(50)

OpenStudy (goldps):

thx, that clears up alot

OpenStudy (alekos):

that's it! well done